Find the tension in cable ?? using the values for ?1, ?2, ?, ? and ???? given in the table. The tension in cable ?? is known to be approximately 97.2236 N.Find the tension in cable AC using the values for L1, L2, a, ß and Mass given in the table. The tension in cable BC is known
Answer #1

Pago ON Resolving the forces TI Cosa sx T2 cos

Here,

Let tension on AC = T1

Tension on BC = T2

We will resolve the tension force into its horizontal and vertical component.

Vertical Component for cable AC

T1sina

= T1sin25

Vertical component of cable BC

=T2sinB

=T2sin41

Also horizontal component for Cable AC

=T1cosa

=T1cos25

Horizontal component of cable BC

=T2cos41

Now if resolve the forces horizontally , we see only two force act and in opposite directions

T1cos25 - T2cos41 =0

T1cos25 =T2cos41

0.90631T1=0.75471 x T2

T1=(0.75471 T2)/0.90631

T1= 0.83273T2

Now forces acting on vertical direction

mg -T1sin25 -T2sin41 = 0

10 x 9.8 - 0.42262T1 -0.65606T2 =0

98 - ( 0.42262 x 0.83273T2) - 0.65606T2 =0

98 - 0.35192T2 -0.65606T2 = 0

1.00798T2 = 98

T2 = 98/1.00798

T2 ≈ 97.2234N ANSWER